Ask your own question, for FREE!
Mathematics 16 Online
OpenStudy (anonymous):

A cell phone company buys computer chips from a supplier. The supplier sends a shipment containing 3% defective chips. Each cell phone uses 3 chips selected independently. What is the probability that all 3 chips will work properly? thankss

OpenStudy (anonymous):

just kidding The probability of selecting a good chip is 97% = .97 Disallowing for the slight change in percentages when a chip is selected, the chances of selecting 3 good ones in a row is: .97 * .97 * .97 = .9127 = 91.27%

OpenStudy (anonymous):

yay me

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!